OpenStudy (anonymous):

hey what is 75/6 just simplify

jimthompson5910 (jim_thompson5910):

find what factors they have in common then divide each piece by that common factor

OpenStudy (anonymous):

omg thank you so much

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

kk

jimthompson5910 (jim_thompson5910):

in this case, 75 and 6 have 3 in common 75/3 = 25 6/3 = 2 so 75/6 = 25/2